OpenStudy (kingj):

What is the 12th term of the arithmetic sequence where a1 = 5 and a6 = 20?

OpenStudy (anonymous):

Arithmetic sequence implies delta between terms is constant. In this case, that constant will be (20-5)/5 = 3. So, the 20th term will be a1 + 19*3 = 5 + 57 = 62.